Suppose we are thinking about replacing an old computer with a new one. The old one cost us $1,740,000; the new one will cost, $2,105,000. The new machine will be depreciated straight-line to zero over its five-year life. It will probably be worth about $495,000 after five years. The old computer is being depreciated at a rate of $384,000 per year. It will be completely written off in three years. If we don’t replace it now, we will have to replace it in two years. We can sell it now for $585,000; in two years, it will probably be worth $171,000. The new machine will save us $365,000 per year in operating costs. The tax rate is 24 percent, and the discount rate is 9 percent. a-1. Calculate the EAC for the the old computer and the new computer. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) a-2. What is the NPV of the decision to replace the computer now? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Α. B C D E F G H 1 Particulars 2 1 2 3 4 5 -2105000 365000 87600 277400 101040 378440 365000 87600 277400 101040 378440 36500A B 1 Particulars 2 3 Initial investment -2105000 4 Savings 365000 365000 365000 365000 365000 5 Less: Tax @ 24% =C4*24% =D4*

A 1 Particulars 2. 3 Initial investment 4 Operating cash flows 5 Add: Salvage value 6 Net cash flow 7 FVIF @9% 8 PV of cash fB Α 1 Particulars Year 2 =-(586000+((384000*3)-586000)*24%) 3 Initial investment Operating cash flows 5 Add: Salvage value 6~ m A B C D 1 Year New machine Old machine Difference PVAF @9% PV of Cash Flows 2 0 -2105000 -721840 -1383160 1 -1383160 3784A B D E 1 Year New machine old machine Difference PVAF @9% 20 -2105000 -721840 =B2-C2 =1/(1+9%)^A2 3 1 378440 92160 =B3-c3 =1